# Efficient Computation of Voronoi Diagrams Using Point-in-Cell Tests

### v1

@@ Abstract.
Since the Voronoi diagram appears in many applications, the topic of improving its computational efficiency remains attractive. We propose a novel yet efficient method to compute Voronoi diagrams bounded by a given domain, i.e., the clipped or restricted Voronoi diagrams. The intersection of the domain and a Voronoi cell (domain-cell intersection) is generated by removing the part outside the cell from the domain, which can be accomplished by several clippings. Different from the existing methods, we present an edge-based search scheme to find clipping planes (bisectors). A test called point-in-cell is first set up to tell whether a space point is in a target Voronoi cell or not. Then, for each edge of the intermediate domain-cell intersection, we will launch a clipping only if its two endpoints are respectively inside and outside the corresponding Voronoi cell, where the bisector for the clipping can be found by using a few times of point-in-cell tests. Therefore, our method only involves the clippings that contribute to the final results, which is a great advantage over the state-of-the-art methods. Additionally, because each domain-cell intersection can be generated independently, we extend the proposed method to the GPUs for computing Voronoi diagrams in parallel. The experimental results show the best performance of our method compared to state-of-the-art ones, regardless of site distribution. This paper was first submitted to SIGGRAPH Asia 2025.
Voronoi diagrams, point-in-cell tests, GPUs
This work is supported by the National Science Foundation of China, under grant 62562046, grant 62272402 and grant 62372389. Corresponding author: Zhonggui Chen. Authors’ addresses: Yanyang Xiao and Yao Li, Nanchang University, 999 Xuefu Avenue, Nanchang, Jiangxi 330031, China. Juan Cao and Zhonggui Chen, Xiamen University, 422 South Siming Road, Xiamen, Fujian 361005, China.
journal: TOGjournalvolume: xxjournalnumber: yyarticle: zzjournalyear: 2025publicationmonth: 9copyright: nonedoi: 0000001.0000001_2ccs: Theory of computation Computational geometryccs: Computing methodologies Parallel algorithms
Figure 1. 1M 3D Voronoi cells bounded by a cube. The results are generated by running our method in parallel on an NVidia RTX 3080 GPU. The site distributions are random (left) and uniform (right). To compute 1M cells, the time costs of the kNN-based method [Ray et al., 2018] and our method are 0.656 vs. 0.155 seconds (left) and 0.063 vs. 0.052 seconds (right), respectively. In addition, our method guarantees to produce exact results, while the kNN-based method has no such guarantee. Only a subset of the cells are rendered, each shrunk for visualization.
1. Introduction
@@
Given a set of points (called sites or generators), a Voronoi diagram [Aurenhammer, 1991] decomposes the space into small cells (called Voronoi cells). Each cell consists of points whose distance to the corresponding site is not greater than that to the other sites, where the distance metric is usually the Euclidean distance. The Voronoi diagram is frequently used in a wide range of applications, including mesh generation [Du et al., 1999; Yan et al., 2013; Jie et al., 2014; Cao et al., 2022], surface remeshing [Yan et al., 2009; Lévy and Liu, 2010; Lévy and Bonneel, 2013; Nivoliers et al., 2015], sampling [Balzer et al., 2009; Chen et al., 2012; Yan et al., 2015], image processing [Chen et al., 2014; Duan and Lafarge, 2015; Cao et al., 2018], point cloud processing [Boltcheva and Lévy, 2017; Chen et al., 2018; Xiao et al., 2025], etc. Improving the computational efficiency of Voronoi diagrams remains attractive, especially for those applications requiring iterative construction of Voronoi diagrams.
@@
For ordinary Voronoi diagrams, some Voronoi cells are unbounded, containing infinitely distant points in the space. In practice, given a domain composed of a set of triangles/tetrahedra, most applications are interested in the intersections of the Voronoi cells and the domain (domain-cell intersections), which are called clipped or restricted Voronoi cells. Thus, our goal is to compute domain-cell intersections as fast as possible, each of which consists of triangle/tetrahedron-cell intersections.
@@
The computation of Voronoi diagrams has been well studied in recent decades. Although many methods have been presented, each of them is with inherent defects. Some algorithms including divide-and-conquer [Shamos and Hoey, 1975], incremental insertion [Green and Sibson, 1978], and sweep-line [Fortune, 1987] were developed only for the planar domains and are hard to use in higher dimensional spaces. Despite jump flooding algorithm (JFA)-based method [Rong and Tan, 2006; Rong et al., 2011] being capable of dealing with 2D and 3D cases, it was designed to cluster pixels from a preset texture for parallel computation on GPUs, thus generating approximate Voronoi cells and being incompatible with irregular domains. To obtain Voronoi diagrams of general cases, the Delaunay triangulation (DT)-based [Aurenhammer, 1991; Yan et al., 2009, 2013] and k-nearest neighbors (kNN)-based methods [Lévy and Bonneel, 2013; Ray et al., 2018; Liu et al., 2022] are more popular. However, the former highly relies on the Delaunay triangulation of the given sites, while the construction of triangulation is time-consuming. The latter which solely depends on kNN queries has attracted much attention in recent years. The cell generation is independent of each other and suitable for parallel computation [Han et al., 2017; Ray et al., 2018; Liu et al., 2022]. The main drawback of kNN-based method is that its efficiency is sensitive to the distribution of the given sites [Ray et al., 2018]. In high dimensional spaces, the kNN-based method usually performs worse than the DT-based method when the distribution leads to long and thin Voronoi cells. Another method called corner validation [Sainlot et al., 2017] shows better performance than the kNN-based method when the sites are far away from the domain. However, if a triangle/tetrahedron intersects with an excess of Voronoi cells, the corner validation method has to run more iterations on that triangle/tetrahedron, resulting in abundant computations.
@@
We propose an efficient method to compute clipped or restricted Voronoi diagrams, where the generation of triangle/tetrahedron-cell intersections is fundamentally distinct from existing methods. We first set up a point-in-cell test to tell whether a space point is in a target Voronoi cell or not. When computing a triangle/tetrahedron-cell intersection, at most one clipping will be launched for each edge of the intersection, where the bisector for the clipping can be found by using a few times of the point-in-cell tests. Thus, our method only involves valid clippings to the final results, which is a great advantage over the existing methods, including DT-based, kNN-based, and corner validation, since a part of their clippings are useless to results but will still be executed. This is the key to our method in obtaining higher efficiency than existing methods. The main contributions of this paper can be summarized as:
@@
  • A novel solution for generating triangle/tetrahedron-cell intersections is presented. For each edge of an intersection, we launch at most one clipping that is guaranteed-valid to the result, where the corresponding bisector can be found quickly by using the point-in-cell tests.
@@
  • By modifying the presented solution for triangle/tetrahedron-cell intersection generation, the proposed method can be easily adapted to GPUs for computing clipped or restricted Voronoi diagrams in parallel. Our parallel algorithm is demonstrated to achieve higher efficiency compared to the state-of-the-art kNN-based algorithm.
@@
In the rest of this paper, section 2 reviews related work on the computation of Voronoi diagrams, and section 3 provides the necessary preliminaries. Our proposed method is introduced in section 4, followed by its parallel implementation in section 5. Implementation details and experimental results are presented in sections 6 and 7, respectively. Finally, section 8 concludes the paper.
2. Related Work
@@
We focus on methods capable of computing both 2D and 3D Voronoi cells within a given domain. A triangle/tetrahedron-cell intersection is typically generated by some clippings, while finding the corresponding bisectors varies across different methods. A clipping is contributed or valid only if it influences the correctness of the final result. Our method only involves valid clippings, in contrast to state-of-the-art approaches such as DT-based and kNN-based methods, which may incorporate invalid clippings.
2.1. DT-based method
@@
It is known that Voronoi diagram is the dual graph of Delaunay triangulation, based on which, the Voronoi diagram can be computed by using its duality [Aurenhammer, 1991; Yan et al., 2009, 2013]. Once the Delaunay triangulation of the given sites is established, we can explicitly obtain its incident neighbors, so that the corresponding Voronoi cell can be generated by clipping the domain using the bisectors of the site and its neighbors. The key to this method is the construction of Delaunay triangulation. Although there have been many algorithms, including flipping [Lawson, 1972; Joe, 1989], incremental insertion [Bowyer, 1981; Watson, 1981; Guibas and Stolfi, 1985; Guibas et al., 1992], and divide-and-conquer [Dwyer, 1987; Cignoni et al., 1998], the time cost to build triangulation is still high for Voronoi generation.
@@
Another issue of the DT-based method being easily ignored is the existence of invalid clippings during the generation of domain-cell intersections. Imagine the case that a Voronoi cell covers multiple triangles/tetrahedra. To generate the triangle/tetrahedron-cell intersection, only a part of the clippings related to the Voronoi cell are useful, while the rest are useless but will still be executed, causing efficiency loss. See Fig. 2(b) as an example, only two clippings are valid for generating the intersection (light green), with the relevant bisectors originating from the red site and two purple sites.
@@
The convex hull-based method [Brown, 1979] lifts the given sites to the space of one dimension higher and constructs a convex hull of these lifted points for computing Voronoi diagrams. It is essentially the same as the DT-based method, because the connectivity of the convex hull is identical to that of the Delaunay triangulation, thereby exhibiting similar drawbacks.
2.2. kNN-based method
@@
Each pair of sites divides the space into two halves, and all the half-spaces associated with a site intersect in a common area, namely its Voronoi cell. To obtain it, we can construct the bisectors of the site and all the others to clip the given domain. Apparently, it is unnecessary to consider all other sites, as a Voronoi cell is primarily defined by bisectors of its nearest neighbors. Therefore, the kNN-based method [Lévy and Bonneel, 2013] finds k-nearest neighbors of the site, where the neighbors are sorted in ascending order according to their distances to the site, then the bisectors of the site and its neighbors are sequentially used to clip the domain for generating the Voronoi cell. The cell generation completes once the security radius is found, i.e., the maximum distance from the site to its cell is smaller than the distance from the site to any of the rest bisectors. Benefiting from the independence of cell generation, the parallelization of this method on GPUs has been well studied [Han et al., 2017; Ray et al., 2018; Liu et al., 2022]. Moreover, this method can also be extended to compute power diagrams [de Goes et al., 2015; Basselin et al., 2021; Xiao et al., 2023].
@@
The kNN-based method suffers from invalid clippings as well. If a site dominates points far away from the site, the method has to visit more neighbors to clip the domain, since the security radius of such a case is hard to meet. All these clippings will be executed even though only a few of them are valid, explaining that the method is time-consuming once the given sites are not uniform. The example in Fig. 2(b) shows that bisectors of the red site and gray sites are useless to generate the triangle/tetrahedron-cell intersection (light green) if the kNN-based method is applied. Although users can set a maximum number of neighbors to avoid excessive clippings, the resultant cells are possibly an approximation of the exact Voronoi cells if the number is not large enough.
2.3. Parallel algorithms
@@
The parallel computation of Voronoi diagrams on GPUs has remained an active research focus. Given their duality, the parallel construction of Delaunay meshes on GPUs is likewise an area of ongoing interest [Cao et al., 2014]. However, data synchronization between threads is required, resulting in limited parallelism for Delaunay mesh construction. In contrast, the JFA-based method [Rong and Tan, 2006; Rong et al., 2011] processes a GPU texture by identifying the nearest site for each pixel, generating approximate Voronoi diagrams even when all pixels are correctly labeled. As mentioned above, parallel implementations of the kNN-based method have been proposed [Han et al., 2017; Ray et al., 2018; Liu et al., 2022]. Specifically, the algorithm of Ray et al. [Ray et al., 2018] works on regular domains, while those of Han et al. [Han et al., 2017] and Liu et al. [Liu et al., 2022] are designed for surface meshes and tetrahedral meshes, respectively. To leverage the GPU architecture, these algorithms typically assign a fixed number of neighbors to each site. However, some cells may require additional neighbors to meet the security radius. This inherent constraint explains why kNN-based algorithms cannot guarantee exact results. As an advantage, our parallelized implementation ensures exact results via the proposed novel generation of triangle/tetrahedron-cell intersection. In addition, for irregular domains, failure to identify all Voronoi cells intersecting with a triangle/tetrahedron can also introduce inaccuracies [Liu et al., 2022], which can be avoided by using a propagation approach [Han et al., 2017].
3. Preliminaries
3.1. Clipped or restricted Voronoi diagram
@@
(1) $\Omega_{i}=\left\{\mathbf{x}\in\mathbb{R}^{d}|\|\mathbf{x}-\mathbf{v}_{i}\|\leq\|\mathbf{x}-\mathbf{v}_{j}\|,\forall j\neq i\right\},i,\dots,n.$
@@
Suppose that the given domain $\mathcal{M}$ consists of a set of small triangles/tetrahedra, i.e., $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$, our goal is to compute the intersections $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$ of the Voronoi cells and the domain, i.e., each site $\mathbf{v}_{i}$ dominates a sub-domain $\Omega_{i|\mathcal{M}}=\Omega_{i}\bigcap\mathcal{M},i=1,\dots,n$. The set of domain-cell intersections $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$ provides a partition of the domain, it is called clipped Voronoi diagram if $\mathcal{M}$ is volumetric, or restricted Voronoi diagram if $\mathcal{M}$ is a surface mesh.
3.2. Edge/face flags of polygon/polytope

### v2

@@ Abstract
Since the Voronoi diagram appears in many applications, the topic of improving its computational efficiency remains attractive. We propose a novel yet efficient method to compute Voronoi diagrams bounded by a given domain, i.e., the clipped or restricted Voronoi diagrams. The intersection of the domain and a Voronoi cell (domain-cell intersection) is generated by removing the part outside the cell from the domain, which can be accomplished by several clippings. Different from the existing methods, we present an edge-based search scheme to find clipping planes (bisectors). A test called point-in-cell is first set up to tell whether a space point is in a target Voronoi cell or not. Then, for each edge of the intermediate domain-cell intersection, we will launch a clipping only if its two endpoints are respectively inside and outside the corresponding Voronoi cell, where the bisector for the clipping can be found by using a few times of point-in-cell tests. Therefore, our method only involves the clippings that contribute to the final results, which is a great advantage over the state-of-the-art methods. Additionally, because each domain-cell intersection can be generated independently, we extend the proposed method to the GPUs for computing Voronoi diagrams in parallel. The experimental results show the best performance of our method compared to state-of-the-art ones, regardless of site distribution. This paper was submitted to IEEE Transactions on Visualization and Computer Graphics in February 2026.
I Introduction
@@
Given a set of points (called sites or generators), a Voronoi diagram (VD) [2] decomposes the space into small cells (called Voronoi cells). Each cell consists of points whose distance to the corresponding site is not greater than that to the other sites, where the distance metric is usually the Euclidean distance. The Voronoi diagram is frequently used in a wide range of applications, including mesh generation [17, 47, 27, 9], surface remeshing [46, 31, 30, 35], sampling [3, 13, 45], image processing [12, 18, 8], point cloud processing [5, 14, 44], etc. Improving the computational efficiency of Voronoi diagrams remains attractive, especially for those applications requiring iterative construction of Voronoi diagrams.
@@
For ordinary Voronoi diagrams, some Voronoi cells are unbounded, containing infinitely distant points in the space. In practice, given a domain composed of a set of simplices (triangles in the 2D and surface cases or tetrahedra in the 3D case), most applications are interested in the intersections of the Voronoi cells and the domain (domain-cell intersections), which are called clipped or restricted Voronoi cells. Thus, our goal is to compute domain-cell intersections as fast as possible, each of which consists of simplex-cell intersections.
@@
The computation of Voronoi diagrams has been well studied in recent decades. Although many methods have been presented, each of them is with inherent defects. Some algorithms including divide-and-conquer [40], incremental insertion [23], and sweep-line [21] were developed only for the planar domains and are hard to use in higher dimensional spaces. Despite jump flooding algorithm (JFA)-based method [38, 37] being capable of dealing with 2D and 3D cases, it was designed to cluster pixels from a preset texture for parallel computation on GPUs, thus generating approximate Voronoi cells and being incompatible with complex domains. To obtain Voronoi diagrams of general cases, the Delaunay triangulation (DT)-based [2, 46, 47] and k-nearest neighbors (kNN)-based methods [30] are more popular. However, the former highly relies on the Delaunay triangulation of the given sites, while the construction of triangulation is time-consuming. The latter which solely depends on kNN queries has attracted much attention in recent years. The cell computation is independent of each other and suitable for parallel computation [36, 26, 32, 4]. The main drawback of kNN-based method is that its efficiency is sensitive to the distribution of the given sites [36]. In high dimensional spaces, the kNN-based method usually performs worse than the DT-based method when the distribution leads to long and thin Voronoi cells. Another method called corner validation [39] shows better performance than the kNN-based method when the sites are far away from the domain. However, if a simplex intersects with a large number of Voronoi cells, the corner validation method requires more iterations, resulting in a significant computational cost.
(a) white noise
(b) blue noise
Figure 1: Computing 5M 3D Voronoi cells bounded by a cube, where the given sites are white noise (a) and blue noise (b), respectively. Tested on an NVIDIA RTX 3080 GPU, the runtimes of two parallel implementations [36, 4] of the kNN-based method and our method are (a): 3.12, 2.81, 0.49 seconds and (b): 0.34, 0.42, 0.26 seconds, respectively. Only a subset of the cells are rendered, each shrunk for visualization.
@@
We propose an efficient method to compute clipped or restricted Voronoi diagrams, where the computation of simplex-cell intersections is fundamentally distinct from existing methods. We first set up a point-in-cell test to tell whether a space point is in a target Voronoi cell or not. When computing a simplex-cell intersection, at most one clipping will be launched for each edge of the intersection, where the bisector for the clipping can be found by using a few times of the point-in-cell tests. Thus, our method only involves valid clippings to the final results, which is a great advantage over the existing methods, including DT-based, kNN-based, and corner validation, since a part of their clippings are useless to results but will still be executed. This is the key to our method in obtaining higher efficiency than existing methods. The main contributions of this paper can be summarized as:
@@
  • A novel solution for generating simplex-cell intersections is presented. For each edge of an intersection, we launch at most one clipping that is guaranteed-valid to the result, where the corresponding bisector can be found quickly by using the point-in-cell tests.
@@
  • By modifying the presented solution for simplex-cell intersection computation, the proposed method can be easily adapted to GPUs for computing clipped or restricted Voronoi diagrams in parallel. Our parallel algorithm is demonstrated to achieve higher efficiency compared to the state-of-the-art kNN-based algorithm.
@@
In the rest of this paper, Section II reviews related work on the computation of Voronoi diagrams, and Section III provides the necessary preliminaries. Our proposed method is introduced in Section IV, followed by its parallel implementation in Section V. Implementation details and experimental results are presented in Sections VI and VII, respectively. Finally, Section VIII concludes the paper.
II Related Work
@@
We focus on methods capable of computing both 2D and 3D Voronoi cells within a given domain. A simplex-cell intersection is typically generated by some clippings, while finding the corresponding bisectors varies across different methods. A clipping is contributed or valid only if it influences the correctness of the final result. We propose a novel solution for generating simplex-cell intersections that involves only valid clippings. This contrasts with state-of-the-art approaches, such as DT-based and kNN-based methods, which may incorporate invalid ones.
II-A DT-based method
@@
It is known that Voronoi diagram is the dual graph of Delaunay triangulation, based on which, the Voronoi diagram can be computed by using its duality [2, 46, 47]. Once the Delaunay triangulation of the given sites is established, we can explicitly obtain incident neighbors of each site, so that the corresponding Voronoi cell can be generated by clipping the domain using the bisectors of the site and its neighbors. The key to this method is the construction of Delaunay triangulation. Although there have been many algorithms, including flipping [29, 28], incremental insertion [6, 42, 25, 24], and divide-and-conquer [19, 15], the time cost to build triangulation is still high for Voronoi computation.
@@
Another issue of the DT-based method being easily ignored is the existence of invalid clippings during the computation of domain-cell intersections. Imagine the case that a Voronoi cell covers multiple simplices. To generate the simplex-cell intersection, only a part of the clippings related to the Voronoi cell are useful, while the rest are useless but will still be executed, causing efficiency loss. See Fig. 2(b) as an example, only two clippings are valid for generating the intersection (light green), with the relevant bisectors originating from the red site and two purple sites.
@@
The convex hull-based method [7] lifts the given sites to the space of one dimension higher and constructs a convex hull of these lifted points for computing Voronoi diagrams. It is essentially the same as the DT-based method, because the connectivity of the convex hull is identical to that of the Delaunay triangulation, thereby exhibiting similar drawbacks.
II-B kNN-based method
@@
Each pair of sites divides the space into two halves, and all the half-spaces associated with a site intersect in a common area, namely its Voronoi cell. To obtain it, we can construct the bisectors of the site and all the others to clip the given domain. Apparently, it is unnecessary to consider all other sites, as a Voronoi cell is primarily defined by bisectors of its nearest neighbors. Therefore, the kNN-based method [30] finds k-nearest neighbors of the site, where the neighbors are sorted in ascending order according to their distances to the site, then the bisectors of the site and its neighbors are sequentially used to clip the domain for generating the Voronoi cell. The cell computation completes once the security radius is found, i.e., the maximum distance from the site to its cell is smaller than the distance from the site to any of the rest bisectors. Benefiting from the independence of cell computation, the parallelization of this method on GPUs has been well studied [36, 26, 32, 4]. Moreover, this method can also be extended to compute power diagrams [4, 16, 43].
@@
The kNN-based method suffers from invalid clippings as well. If a site dominates points far away from the site, the method has to visit more neighbors to clip the domain, since the security radius of such a case is hard to meet. All these clippings will be executed even though only a few of them are valid, explaining that the method is time-consuming once the given site distribution does not exhibit blue noise properties. The example in Fig. 2(b) shows that bisectors of the red site and gray sites are useless to generate the simplex-cell intersection (light green) if the kNN-based method is applied. Although users can set a maximum number of neighbors to avoid excessive clippings, the resultant cells are possibly an approximation of the exact Voronoi cells if the number is not large enough.
II-C Parallel algorithms
@@
The parallel computation of Voronoi diagrams on GPUs and multi-cores CPUs has remained an active research focus. Given their duality, the parallel construction of Delaunay triangulations is likewise an area of ongoing interest (e.g., [10, 34, 11, 22]). Compared to traditional algorithms (e.g., [10, 34]) that require data synchronization, the methods based on the local Delaunay lemma [11, 22] independently construct the one-ring neighboring Delaunay simplices for each site, thus enabling a higher degree of parallelism. Since the intersections of the given domain and Voronoi cells are central to computing Voronoi diagrams, we focus on parallel algorithms for directly generating clipped (or restricted) Voronoi cells. The JFA-based method [38, 37] processes a GPU texture by identifying the nearest site for each pixel, generating approximate Voronoi diagrams even when all pixels are correctly labeled. As mentioned above, parallel implementations of the kNN-based method have been proposed [36, 26, 32, 4]. Specifically, the algorithm of Ray et al. [36] works on simple domains, while those of Han et al. [26] and Liu et al. [32] are designed for surface meshes and tetrahedral meshes, respectively. To leverage the GPU architecture, these algorithms typically assign a fixed and sufficiently large number of neighbors to each site, imposing a high memory overhead on the GPU. To reduce memory usage, Basselin et al. [4] presented a multi-pass parallel scheme, starting with a small number of neighbors and increasing it iteratively for Voronoi cells that do not meet the security radius in subsequent passes.
Our method can be easily extended to GPUs. The procedure for generating simplex-cell intersections is adapted to allow independent computation, thereby enabling a parallel algorithm that operates without inter-thread data synchronization on GPUs. Compared to the aforementioned kNN-based implementations, our parallel algorithm not only uses less memory but also delivers faster results.
III Preliminaries
III-A Clipped or restricted Voronoi diagram
@@
$\Omega_{i}=\left\{\mathbf{x}\in\mathbb{R}^{d}|\|\mathbf{x}-\mathbf{v}_{i}\|\leq\|\mathbf{x}-\mathbf{v}_{j}\|,\forall j\neq i\right\},i,\dots,n.$ (1)
@@
Suppose that the given domain $\mathcal{M}$ consists of a set of small simplices, i.e., $\mathcal{M}=\left\{\mathcal{T}_{a}\right\}_{a=1}^{m}$, our goal is to compute the intersections $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$ of the Voronoi cells and the domain, i.e., each site $\mathbf{v}_{i}$ dominates a sub-domain $\Omega_{i|\mathcal{M}}=\Omega_{i}\bigcap\mathcal{M},i=1,\dots,n$. The set of domain-cell intersections $\left\{\Omega_{i|\mathcal{M}}\right\}_{i=1}^{n}$ provides a partition of the domain, it is called clipped Voronoi diagram if $\mathcal{M}$ is volumetric, or restricted Voronoi diagram if $\mathcal{M}$ is a surface mesh.
III-B Edge/face flags of simplex-cell intersection